VIS Meeting on 2016/5/20
2016/5/20 13:30 -
Participants:Sato, Hirata, Ishizaki, Mark, Fabian, Takahashi, Ohishi, Aso, Shoda, Raffaele, Okutomi, Akutsu
The current overall schedule chart.
iKAGRA related activities
- PR3 post-run work status (Shoda)
IP tests in Kamioka
- #Control test may be performed with the hanging test of BS
- Ask Joris what is his intention and schedule (Raffaele)
- Ask if we can perform the control test remotely (Aso)
- Suspended bread board
- #Many critical components will be installed on it (baffles, steering mirrors for the green injection etc)
- We have to make sure that the bread board suspension is properly damped
- Takahashi checked Sekiguchi thesis
- Damping effect depends on the ratio of mass and resonant frequency. Estimated effect was 0.45 for the experiments. The number of magnets needed in the same condition is about half.
30 magnets necessary for the gap of 5mm -> need to update the calculation for the double pendulum case (Takahashi)
- Accelerometers for IP
Geophone is too noisy for innertial damping in Kamioka -> this argument presented in Sekiguchi PhD thesis
- #TAMA SAS accelerometers for Type-A
- NIKHEF accelerometers for BS
- Joris seems to be interested in continue the test according to Raffaele.
- Need to see if this accelerometer is good enough (seems to have some problems).
- #Which accelerometers for SR ?
- Buy from NIKHEF ? How much is the cost ?
- Get the design from NIKHEF and produce them in Japan (ATC) ?
Conclusions of this week:
We should first see the result of Fujii-kun's test. Aso will follow up on his experiment in the next week.
For the question of how bad the geophone is, Shoda will check Sekiguchi thesis and report by email.
- Pre-isolator assembly and tests
- Sato-san will come to Kamioka next week
- What is the plan for SR Type-B
Takahashi and Mark will start the discussion on, where, who, when -> continue.
- Bottom filter local sensor for Type-A
Tell SEO about the wider cross tube -> Wider tube is not possible
Fix EQ stop at the bottom ? -> This is possible
- Takahashi had a discussion with Saito: modify the design of the bottom plate is easier.
- #EQ stop design should take into account of the sensor/actuator mounts.
- #Use the same sensor/actuator as Type-Bp ?
- #Sato-san is responsible for this part.
- Dummy payload design for Type-A
- Okutomi-kun is assigned.
Okutomi will make schematic design. One of the engineers will make drawings. Report more detailed plan by the next week (Takahashi).
- Wide OSEM production (Akutsu)
- For BS test hanging, we need the wider version by mid-to-late June.
- Got cost estimates for the small parts. Will order for 20 OSEMs.
Length sensing OpLev
- Discuss in Optics Meeting.
- #In order to test this with the BS test haning, we want to have wide OSEMs as witness sensors on the mirror.
- #Stab-flag and tipped flag are both in quotation. We will order both. The gluing jigs will be the same.
#Who will design and prepare the BS length sensing OpLev ? -> Ask Simon (Aso) -> in progress
- #What shall we do with PR2
PR2 chamber has only one viewport avaibale at 45deg. angle for OpLev
Original idea: Put a reflecting mirror on the bread board and monitor the rotation of the bread board with another OpLev.
- No length sensing
Two OpLevs needed
Another idea: Introduce the light with a fiber into the vacuum chamber to form a one-way 45deg. OpLev for length sensing.
- Length sensing possible
- Affected by the vacuum chamber tilt (if we mount the fiber collimator to the inner frame, the tilt will be less).
- Need to find a place to mount the fiber collimator inside the vacuum chamber
See if we can find any mounting point on the inner frame (Akutsu) by the next next VIS meeting.
- What are necessary in the next few weeks ?
- We need to count the number of necessary circuits for Type-A and Type-Bp, so that we can tell AEL the total number necessary for VIS.
Make a list by the end of May (Shoda, Takahashi) -> List
- Presented the list to AEL. Miyakawa-san agreed.
- Need to update the list to include IMMT
- Discuss how many spares to make at the next AEL meeting
- Watch dog system (Tanaka)
- Wants to do tests with the existing PR3 sometime in fall ? Need coordination.
- photo-MOS relay switch resistance fluctuation
Noise requirement for the resistance fluctuation is necessary (Aso)
Tanaka measured the photo-MOS relay resistance fluctuation-> 37mOhm RMS upper limit
- Will make spectrum of the resistance fluctuation.
DC-DC convertor noise check: quiter than a battery ?
#The next step: Make a prototype with slow shutdown mechanism
- Past Week Summary
- Fabian did inventory of parts/fasteners for dummy BS, RM, IM and IRM - almost all at Mitaka. A handful (picos, pico drivers) will be brought back from Kamioka.
- Fabian continued work on Mathematica and Matlab models of the BS and control system. He now gets much better agreement after correcting discrepancies between the parameter sets.
- Hirata-san worked on:
- Checking Mirapro assembly frame drawings
- Ordering missing screws and small items (screws for dummy mass, fine pitch M8 for RM)
- Work platforms and clean booth second floor arrangement.
Load gauge adapter and associated parts
- Possible fall-back plans for the rods connecting the security structure to the PI.
- Design for additional IM trim masses (existing design had one mass per corner and no provision for adding/removing).
Mark designed a guide part to help place the PI accurately onto the security structure.
Some holes are missing on the bottom of the BF: 4 holes instead of 6 -> less number of the cable cramps
- Put multiple cables in one cable cramp.
- Check the SF too.
- IM assembly
- Planning to do it in ATC. Use the clean room used for the SF assembly.
- Fabian made an inventory. Most parts are ready.
- Can we make SR wire breakers bigger?
Fukushima-san preparing test: 3/4/2016 + "1 month" = ≈4/1/2016, not ready as of 4/8.-> Takahashi will ask the status.
- Flags, Metal Prisms
- IM and BS flags and metal prisms ordered; expected 5/31
- Sapphire Prisms - expected 6/5
- BS dummy optic RECEIVED
- Mirror box (Tatsumi)
- Tatsumi-san has ordered modifications to mirror box for final flags and prisms.
- An extra structure attached to the EQ stop to prevent tipping over of the BS.
- Will consider if we can make it to the test hanging. But not critical for the test hanging.
- No action since last week.
- Not enough crane height problem (Hirata, Saito)
- No action since last week
- EQ stop screws, extensions and lock nuts (in manufacture - promised for 4/22)
- Ballast masses for payload
- Initial batch received - need to design and order smaller ones to allow for adjustments
- Lower breadboard stuff
- Breadboard spring bases and associated cable clamps RECEIVED
- Breadboard springs (in manufacture)
- Damper ring mods RECEIVED
- Nail-head parts: quote from Mirapro received, not urgent, will be combined with assembly frame order.
- BS Assembly Frame Design (Mark, Hirata)
- Hirata-san did new Inventor assembly with Mirapro's initial design plus various improvements including clean bag supports.
- Clean bag
- Ordered - expected 6/8.
- 2D and 3D CAD
- BS 3D Assembly
- Load frame adapter added
- Item list and timetable was given to Miyakawa-san - some corrections received and added to schematic but not yet posted.
- Status on AEL boxes OK according to Miyakawa-san.
- Signal generator and cable to be used for BS test. May order NIKHEF cards later if concern about AC-driven units in racks.
- Need to get new longer versions of some previously ordered BF-PI and SF-PI cables (Takahashi).
- Need to check previously ordered TM-BF and IM-BF OSEM cables have been flipped (Takahashi).
- Decided to hand-make cables for geophones because previously ordered cables have no conductor on pin 5.
Stepper drive box (requested by Sekiguchi-san T1503385) is being designed, should be ready in time.
- Procedure documents
Mark worked on BS Assembly Procedure E1504235 draft
Task list (Takahashi)
Clean booth design for 2F https://www.dropbox.com/s/8bi9ij91radmpt5/IXYV160512.pdf?dl=0
- Ask Miyoki-san to add cable racks
- Prototype test of the EQ stop
- Will do the test in ATC
- EQ stop and assembly frame prototype were shipped to Mitaka
- Will start the preparation from the next
- Sensors to monitor the release of the SFs from the EQ stop.
- Now considering to use some touch sensors
- Will test this sensor in the prototype test
- #The test of prototype long maraging rod was finished
- Will order them with maraging rods for the cryogenic payload early July. The lead time is 2 months.
- Type-B Standard Filters preparation
- 6th SF measurements were done ?
- Shipped the SF for BS to Mitaka.
- Ship the measured SFs to Kamioka at an apropriate time
- Type-A SF preparation
Needs to cut more finshing rod blades -> ATC is working on it now.
- The next Type-A meeting will be on 5/26.
- Simultion (Okutomi)
- #Local sensors for BF
- Modification for bKAGRA
- Discussing with Saito-san of ATC
- Specified what can be modified
- Asked to fix the EQ stop to the traverser
- Asked Fujii-kun to work on the noise requirements for the BF sensor/actuator
- Still waiting for the LVDT design from Virgo
- Full CAD model of the original Type-Bp needs to be created
Site work plan for the next week
- Mon: IP
- Tue: IP
- Wed: IP
- Thu: IP
- Fri: IP
- Travel (Week of 5/23)
- Fujii: 23-27 (Accelerometer, PR3)
- Sato: Wed-Fri (Acceleromter)
- Travel (Week of 5/30):
Mark is away: July 8-20
- #GAS blade inspection by Virgo (Takahashi)
- Shipped the blades. Will get results in 1-2 months.
2016/5/27(Fri) 13:30 -